Grade 9Mechanics

If a vector+ b vector=c vector and a^2+=b^2+c^2,then prove that a and b are perpendicular to eCh other

Biswajit debnath
7 Years agoGrade 9

1 Answer

AKSHANSH PAREEK
7 Years ago
Given A+B=C and A^2+B^2=C^2.
(A+B).(A+B)=C.C
=>A.A+A.B+B.A+B.B=C.C
=>A^2+2A.B+B^2=C^2
=>2A.B=0
=>A.B=0
=>|A||B|cosθ=0
=>cosθ=0
=>θ=π/2
 
please  approve it